Marla’s Massages and More bought a special massage table two years ago for $9,300. At the time, the accountants told Marla to depreciate the table using straight line depreciation over three years. Now there is a new and improved massage table out that Marla would like to buy, but she can get only $2,100 for the old table. Marla is in the 35% tax bracket. What is the after tax salvage value of the old table?
Answer #1

Annual depreciation=(Cost-Residual value)/Useful Life

=(9300/3)=$3100/year

Hence book value as on date of sale=Cost-Accumulated Depreciation

=9300-(3100*2)=$3100

Hence loss on sale=3100-2100=$1000

Hence after tax salvage value=Sale proceeds+(Tax rate*Loss on sale)

=2100+(1000*35%)

=$2450